# Indoor Multi-View Radar Object Detection via 3D Bounding Box Diffusion

### v1

@@ Appendix A Details on Diffusion Models
Denoising Diffusion Probabilistic Models (DDPM) (Ho2020_DDPM) and Denoising Diffusion Implicit Models (DDIM) (Song2021_DDIM) are latent variable models designed to approximate the data distribution $q(\boldsymbol{x}_{0})$ using a generative model distribution $p_{\theta}(\boldsymbol{x}_{0})$. The generative model structure is expressed as:
@@
For REXO training in Fig. 4 and inference, we construct a $\mathtt{DenoisingDet}_{\theta}$ that directly predicts $\boldsymbol{x}_{0}$, similar to the above DDIM and DiffusionDet  (Chen2023_diffusiondet). This direct estimation of $\boldsymbol{x}_{0}$ allows us to compute the set-prediction loss in DETR (Carion2020_detr) over both radar and camera coordinate systems. As a result, the noise prediction network $\boldsymbol{\epsilon}^{\left({\tau_{i}}\right)}$ is a function of the predicted $\hat{\boldsymbol{x}}_{0}$.
@@ Appendix B Details of Cross-View Radar-Conditioned Denoising Detector
The BBox detector module, $\mathtt{BBoxDet}_{\theta}$, first incorporates a time-dependent predictor, $\mathtt{Predictor}$, designed to extract time-dependent features $\boldsymbol{f}$ from the aligned cross-view radar features $\boldsymbol{Z}_{\mathtt{radar}}^{\mathtt{crop}}$ and the time-embedding of $t$ (Ho2020_DDPM),
@@ Appendix D RFMask3D as a Baseline
Model
Backbone - ResNet18 ResNet18 ResNet18 ResNet18
@@
Training
Batch size - 32 32 32 32
@@
As one of the baselines in our evaluation experiments, we constructed RFMask3D by extending RFMask (Wu2023_RFMask) to 3D. RFMask uses a region proposal network (RPN) to extract regions of interest (RoIs) from a horizontal heatmap based on 2D anchor boxes and predicts 3D BBoxes in the 3D radar coordinate system by combining them with fixed heights. By designing an RPN that uses 3D anchor boxes, we explicitly extract 3D RoIs from both horizontal and vertical heatmaps, as shown in Fig. 13, enabling the estimation of 3D BBoxes. Unlike RFMask, this method allows for the learning of height as well.
@@ Appendix F Definition of Metrics
We adopt average precision on intersection over union (IoU) (Everingham2010_pascalvoc) as an evaluation metric. IoU is the ratio of the overlap to the union of a predicted BBox $A$ and annotated BBox $B$ as:
@@
Average recall (AR) (Hosang2016_AverageRecall) between 0.5 and 1 of $\mathtt{IoU\;overlap\;threshold}$ can be computed by averaging over the overlaps of each annotation $\mathrm{gt}_{i}$ with the closest matched proposal, that is integrating over the $y:\mathtt{recall}$ axis of the plot instead of the $x:\mathtt{IoU\;overlap\;threshold}$ axis. Let $o$ be the $\mathtt{IoU}$ overlap and $\mathtt{recall}\left({o}\right)$ the function. Let $\mathtt{IoU}\left({\mathrm{gt}_{i}}\right)$ denote the $\mathtt{IoU}$ between the annotation $\mathrm{gt}_{i}$ and the closest detection proposal:

### v2

@@ Appendix A Details on Diffusion Models
Denoising Diffusion Probabilistic Models (DDPM) (Ho et al. 2020) and Denoising Diffusion Implicit Models (DDIM) (Song et al. 2021) are latent variable models designed to approximate the data distribution $q(\boldsymbol{x}_{0})$ using a generative model distribution $p_{\theta}(\boldsymbol{x}_{0})$. The generative model structure is expressed as:
@@
For REXO training in Fig. 4 and inference, we construct a $\mathtt{DenoisingDet}_{\theta}$ that directly predicts $\boldsymbol{x}_{0}$, similar to the above DDIM and DiffusionDet  (Chen et al. 2023b). This direct estimation of $\boldsymbol{x}_{0}$ allows us to compute the set-prediction loss in DETR (Carion et al. 2020) over both radar and camera coordinate systems. As a result, the noise prediction network $\boldsymbol{\epsilon}^{\left({\tau_{i}}\right)}$ is a function of the predicted $\hat{\boldsymbol{x}}_{0}$.
@@ Appendix B Details of Cross-View Radar-Conditioned Denoising Detector
The BBox detector module, $\mathtt{BBoxDet}_{\theta}$, first incorporates a time-dependent predictor, $\mathtt{Predictor}$, designed to extract time-dependent features $\boldsymbol{f}$ from the aligned cross-view radar features $\boldsymbol{Z}_{\mathtt{radar}}^{\mathtt{crop}}$ and the time-embedding of $t$ (Ho et al. 2020),
@@ Appendix D RFMask3D as a Baseline
Model Backbone - ResNet18 ResNet18 ResNet18 ResNet18
@@
Training Batch size - 32 32 32 32
@@
As one of the baselines in our evaluation experiments, we constructed RFMask3D by extending RFMask (Wu et al. 2023) to 3D. RFMask uses a region proposal network (RPN) to extract regions of interest (RoIs) from a horizontal heatmap based on 2D anchor boxes and predicts 3D BBoxes in the 3D radar coordinate system by combining them with fixed heights. By designing an RPN that uses 3D anchor boxes, we explicitly extract 3D RoIs from both horizontal and vertical heatmaps, as shown in Fig. 13, enabling the estimation of 3D BBoxes. Unlike RFMask, this method allows for the learning of height as well.
@@ Appendix F Definition of Metrics
We adopt average precision on intersection over union (IoU) (Everingham et al. 2010) as an evaluation metric. IoU is the ratio of the overlap to the union of a predicted BBox $A$ and annotated BBox $B$ as:
@@
Average recall (AR) (Hosang et al. 2016) between 0.5 and 1 of $\mathtt{IoU\;overlap\;threshold}$ can be computed by averaging over the overlaps of each annotation $\mathrm{gt}_{i}$ with the closest matched proposal, that is integrating over the $y:\mathtt{recall}$ axis of the plot instead of the $x:\mathtt{IoU\;overlap\;threshold}$ axis. Let $o$ be the $\mathtt{IoU}$ overlap and $\mathtt{recall}\left({o}\right)$ the function. Let $\mathtt{IoU}\left({\mathrm{gt}_{i}}\right)$ denote the $\mathtt{IoU}$ between the annotation $\mathrm{gt}_{i}$ and the closest detection proposal: